Circular No. 8462

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA IN M31 M. Della Valle, Istituto Nazionale di Astrofisica (INAF), Arcetri; A.

Pizzella and L. Coccato, Dipartimento di Astronomia, Universitá di Padova; and I. Yegorova, International School for Advanced Studies, Trieste; and E. Giro, INAF and Osservatorio Astronomico di Padova, report: “Preliminary analysis of spectra of the bright variable star discovered recently in M31 (IAUC 8461), obtained on Jan. 10.73, 11.87, and 12.76 with the Asiago Observatory 1.8-m telescope (+ AFOSC; range 400–760 nm, resolution ∼ 0.3 nm), confirms this object to be a classical nova caught near maximum light. The spectrum obtained on Jan. 10 shows a blue, nearly featureless continuum, whereas those taken on Jan. 11 and 12 are characterized by the appearance of faint Hα and Na I emission lines, flanked by P-Cyg profiles, and Fe II lines (mult. 49, 48). The expansion velocity derived from the minima of the P-Cyg profiles are ∼ 900 km/s.”

SUPERNOVA 2005B IN UGC 11066 D. Lane, Saint Mary’s University, reports that he and P. Gray have

discovered an apparent supernova (mag ∼ 18.0) on unfiltered CCD images taken on Jan. 12.002–12.18 UT with a 0.28-m reflector at Stillwater Lake, NS; earlier images from 2004 Dec. 17.02 also show the object near limiting mag ∼ 18.5. SN 2005B is located at α = 17^{h}54^{m}48^{s}.8, δ = +71^{o}32′35′′ (equinox 2000.0), which is ≈ 18′′ west and 10′′ north of the center of UGC 11066. Nothing was visible at this location on earlier images taken on 2004 Aug. 23 and Sept. 15 (limiting mag 19.0). An independent discovery of SN 2005A (at mag 17.5) has been reported by T. Boles (cf. IAUC 8446) on an unfiltered CCD image taken on Jan. 13.227. Boles provides position end figures 48^{s}.72, 34′′.5, adding that nothing is visible at this position on his images from 2004 Oct. 2 and Sept. 4 (limiting mag 19.5), and it is not present on Digitized Sky Survey plates from 1993 (limiting red mag 21.0) and 1989 (limiting blue mag 20.5).

CORRIGENDA On IAUC 8461, ‘Supernovae 2004gq and 2005A’, last line, for (IAUC

8404) read (IAUC 8452)

On IAUC 8460, ‘Comet C/2004 Q2 (Machholz)’, line 5, for 7-mm read 7-cm

On IAUC 8445, ‘Supernova 2004gj in IC 701’, line 5, for eastern read western

2005 January 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green